Pressure

The weight of your 1400 kg car is supported equally by its four tires, each inflated to a gauge pressure of 35.0 lb/in^2.

What is the area of contact each tire makes with the road? (answer must be in cm^2) What gauge pressure is required to give an area of contact of 125 cm^2 for eachtire? (answer must be in lb/in^2)
Answer #1

a) P = F/A

35 = 1400 x 9.8/4A

A = 1400 x 0.4535 x 9.8/4 x 35

A = 44.443 in2

A = 44.443 x 2.542

A = 286.72 cm2

b) P = 1400 x 0.4535 x 9.8/125 x 0.39372

P = 321.137 lb/in^2

answered by: piyush
Answer #2
convert 140 kg to lbs. that 3080 lbs. so 3080 lbs / (35 lbs/in^2) = 88 in^2 total. Then each tire has 88/4 = 22 in^2 of contact.

well, we want each tire to have 125 in^2 of contact? That seems like alot, but sure - just take 125*4 = 500 total inches of contact. now we want to find Pwhere

3080/P = 500 so P = 6.16 lbs/in^2
answered by: Jeniffer